# Fast and Efficient Matching Algorithm with Deadline Instances

### v2

@@ 5 Conclusion
In this paper, we study the online matching problem with $deadline$ and solve it with a sketching matrix. We provided a new way to compute the weight of the edge between two nodes in a bipartite. Compared with original algorithms, our algorithms optimize the time complexity from $O(nd)$ to
$\widetilde{O}(\epsilon^{-2}\cdot(n+d)).$
Furthermore, the total weight of our algorithms is very close to that of the original algorithms respectively, which means the error caused by the sketching matrix is very small. Our algorithms can also be used in areas like recommending short videos to users. For matching nodes with many entries, the experiment result shows that our algorithms only take us a little time if parameter $s$ is small enough. We think generalizing our techniques to other variations of the matching problem could be an interesting future direction. We remark that implementing our algorithm would have carbon release to the environment.
Impact Statement
This paper aims to study the fundamental machine learning problem, namely the online weighted matching problem. There might be some social impacts resulting from this work, but we do not believe it is necessary to highlight them in this paper.
@@ Appendix A Missing Proofs
In Section A.1, we provide the proof for Theorem 3.5. In Section A.2, we provide the proof for Lemma 3.10. In Section A.3, we provide the proof for Lemma 3.11. In Section A.4, we provide the proof for Lemma 3.14. In Section A.5, we provide the proof for Lemma 3.15. In Section A.6, we provide the proof for Lemma 3.16. In Section A.7, we provide the proof for Lemma 3.17.
A.1 Proof of Theorem 3.5
Theorem A.1 (Formal version of Theorem 3.5, Correctness of FastPostponedGreedy in Algorithm 3 and Algorithm 4).
@@
A.2 Proof of Lemma 3.10
Lemma A.2 (Formal version of Lemma 3.10).
@@
A.3 Proof of Lemma 3.11
Lemma A.3 (Formal version of Lemma 3.11, Space storage for FastGreedy in Algorithm 1 and Algorithm 2 ).
@@
A.4 Proof of Lemma 3.14
Lemma A.4 (Formal version of Lemma 3.14).
@@
A.5 Proof of Lemma 3.15
Lemma A.5 (Formal version of Lemma 3.15).
@@
A.6 Proof of Lemma 3.16
Lemma A.6 (Formal version of Lemma 3.16).
@@
A.7 Proof of Lemma 3.17
Lemma A.7 (Formal version of Lemma 3.17, Space storage for FastPostponedGreedy in Algorithm 3 and Algorithm 4).

### v3

@@ 5 Conclusion
In this paper, we study the online matching problem with $deadline$ and solve it with a sketching matrix. We provided a new way to compute the weight of the edge between two nodes in a bipartite. Compared with original algorithms, our algorithms optimize the time complexity from $O(nd)$ to $\widetilde{O}(\epsilon^{-2}\cdot(n+d)).$ Furthermore, the total weight of our algorithms is very close to that of the original algorithms respectively, which means the error caused by the sketching matrix is very small. Our algorithms can also be used in areas like recommending short videos to users.
For matching nodes with many entries, the experiment result shows that our algorithms only take us a little time if parameter $s$ is small enough. We think generalizing our techniques to other variations of the matching problem could be an interesting future direction. We remark that implementing our algorithm would have carbon release to the environment.
@@ Appendix A Missing Proofs
In Section A.1, we provide the proof for Lemma 3.1. In Section A.2, we provide the proof for Lemma 3.2. In Section A.3, we provide the proof for Lemma 3.3. In Section A.4, we provide the proof for Theorem 3.5. In Section A.5, we provide the proof for Lemma 3.10. In Section A.6, we provide the proof for Lemma 3.11. In Section A.7, we provide the proof for Lemma 3.14. In Section A.8, we provide the proof for Lemma 3.15. In Section A.9, we provide the proof for Lemma 3.16. In Section A.10, we provide the proof for Lemma 3.17.
A.1 Proof of Lemma 3.1
Lemma A.1 (Restatement of Lemma 3.1).
StandardGreedy in Algorithm 5 is a $1/2$-competitive algorithm.
Proof.
We use $w_{s}$ to denote the current matching value of seller $s$. Let $w_{s,b}$ denote the weight of the edge between the buyer node and $s$-th seller node. Let $i(b)$ be the increment of matching value when buyer $b$ arrives: $i(b)=\max_{s\in{\cal S}}\{w_{s,b},w_{s}\}-w_{s}.$ Let $v_{f}(s)$ be the final matching value for a seller node $s$. If we add all the increments of matching value that every buyer node brings, we will get the sum of every seller node’s matching value. $\sum_{b\in{\cal B}}i(b)=\sum_{s\in{\cal S}}v_{f}(s).$ Let $\mathrm{alg}$ be the weight of the matching in StandardGreedy. Let $\mathrm{alg}=\sum_{s\in{\cal S}}v_{f}(s)$. Let $\mathrm{opt}$ be the optimal matching weight of any matching that can be constructed. Let $\lambda_{i}$ denote the arrival time of node $i$. We need to do the following things to acquire the maximum of $\mathrm{opt}$:
$\mathrm{minimize}$ $~{}\sum_{i\in[n]}a_{i}$
$\mathrm{subject~{}to}$ $~{}a_{i}\geq 0$ $\forall i\in[n]$
$~{}|\lambda_{i}-\lambda_{j}|\leq\mathrm{dl},$ $\forall i<j$
$~{}a_{i}+a_{j}\geq w_{i,j}$ $\forall i,j\in[n].$
We conclude
A.2 Proof of Lemma 3.2
Lemma A.2 (Restatement of Lemma 3.2).
Let $w_{i,j}$ denote the real weight of the edge between seller $i$ and buyer $j$, and $\widetilde{w}_{i,j}$ denote the approximated weight. Let $\epsilon\in(0,0.1)$ denote the precision parameter. If for all $\epsilon$, there exists an $\alpha$-approximation algorithm for the online weighted matching problem and a $\delta>0$, where $(1-\delta)w_{i,j}\leq\widetilde{w}_{i,j}\leq(1+\delta)w_{i,j}$ for any seller node $i$ and buyer node $j$, there exists a greedy algorithm with competitive ratio $\alpha(1-\epsilon)$.
Proof.
Since $\epsilon\in(0,0.1)$ and we have $(1-\delta)w_{i,j}\leq\widetilde{w}_{i,j}\leq(1+\delta)w_{i,j}$. Let $\mathrm{alg}$ and $\mathrm{opt}$ denote the matching weight and optimal weight of the original problem $G=({\cal S},{\cal B},E)$ respectively. Then we define $\mathrm{alg^{\prime}}$ is the matching weight of another problem $G^{\prime}=({\cal S},{\cal B},E^{\prime})$, and our algorithm gives the optimal weight $\mathrm{opt^{\prime}}$. Since for each edge $\widetilde{w}_{i,j}\leq(1+\delta)w_{i,j}$ and $\mathrm{alg}=\sum_{\begin{subarray}{c}i\in{\cal S}\\ j\in{\cal B}\end{subarray}}w_{i,j}$, we can know that $\mathrm{alg^{\prime}}\leq(1+\delta)\mathrm{alg}.$ Since for each edge $(1-\delta)w_{i,j}\leq\widetilde{w}_{i,j}$ and $\mathrm{opt}=\sum_{\begin{subarray}{c}i\in{\cal S}\\ j\in{\cal B}\end{subarray}}w_{i,j}$, we can know that $\mathrm{opt^{\prime}}\geq(1-\delta)\mathrm{opt}.$ Thus,
$\mathrm{alg}\geq~{}\frac{1}{1+\delta}\mathrm{alg^{\prime}}\geq~{}\frac{\alpha}{1+\delta}\mathrm{opt^{\prime}}\geq~{}\frac{\alpha(1-\delta)}{1+\delta}\mathrm{opt}\geq~{}\alpha(1-2\delta)\mathrm{opt},$
where the first step follows from $\mathrm{alg^{\prime}}\leq(1+\delta)\mathrm{alg}$, the second step follows from there exists a $\alpha$-approximation greedy algorithm, the third step follows from $\mathrm{opt^{\prime}}\geq(1-\delta)\mathrm{opt}$, and the last step follows from $\frac{\alpha(1-\delta)}{1+\delta}\geq\frac{\alpha(1-\delta)^{2}}{1-\delta^{2}}\geq\frac{\alpha(1-\delta)^{2}}{2}\geq\alpha(1-2\delta)$. Then we can draw our conclusion because we can replace $2\delta$ with a new $\epsilon=2\delta$ in the last step. ∎
A.3 Proof of Lemma 3.3
Lemma A.3 (Restatement of Lemma 3.3).
Proof.
Since the probability of a node being a seller is $1/2$, we can know that
A.4 Proof of Theorem 3.5
Theorem A.4 (Restatement of Theorem 3.5, Correctness of FastPostponedGreedy in Algorithm 3 and Algorithm 4).
@@
A.5 Proof of Lemma 3.10
Lemma A.5 (Restatement of Lemma 3.10).
@@
A.6 Proof of Lemma 3.11
Lemma A.6 (Restatement of Lemma 3.11, Space storage for FastGreedy in Algorithm 1 and Algorithm 2 ).
@@
A.7 Proof of Lemma 3.14
Lemma A.7 (Restatement of Lemma 3.14).
@@
A.8 Proof of Lemma 3.15
Lemma A.8 (Restatement of Lemma 3.15).
@@
A.9 Proof of Lemma 3.16
Lemma A.9 (Restatement of Lemma 3.16).
@@
A.10 Proof of Lemma 3.17
Lemma A.10 (Restatement of Lemma 3.17, Space storage for FastPostponedGreedy in Algorithm 3 and Algorithm 4).
@@ Appendix B More Experiments
Appendix C More Related Work
The JL lemma [50] provides strong theoretical guarantees with high probability $(1\pm\epsilon)$-distortion, which is critical for maintaining our competitive ratio proofs. This is evidenced in Lemma 2.5 and its application in Theorems 3.4 and 3.5. Compared with principal component analysis (PCA), sketching typically requires less computational complexity $O(snd)$, where the sketching matrix is $s\times n$ and the data matrix is $n\times d$ and in our case, $s=O(\epsilon^{-2}\log(n/\delta))$ because it requires sampling of rows of the data matrix. However, PCA requires the computation of SVD, which requires $\min\{O(n^{2}d),O(nd^{2})\}$, for an $n\times d$ data matrix. In sketching, the value of $s$ can be adjusted, representing a trade-off between efficiency and accuracy. This trade-off can be also seen in our experimental results (especially in Table 2). Regarding accuracy, sketching can generate more random results because of its sampling procedure and the choice of $s$, whereas PCA has a more deterministic error.
More generally, other theoretical machine learning works study LLM efficiency [96, 115, 26, 18, 109, 108, 66, 74, 94, 72, 106, 73, 67, 71, 92, 60, 54, 22, 19, 59, 65, 25, 62, 55, 29, 58, 27, 14, 17, 114, 61], differential privacy [5, 4, 10, 1, 7, 8, 64], determinantal point processes [104], fast Gaussian transform [47], attack problems [45], kernel density estimation [77], online bipartite matching [45], active learning [30], leverage score [75], reinforcement learning [112, 111, 81, 82, 68, 69], circuit complexity [23, 63], and fairness analysis [24].
Sketching has been applied to many other fields, such as attention approximation [103, 42, 41, 95, 76, 78], $k$ means clustering [70], federated learning [13], tensor decomposition [37], weighted low rank approximation [101], linear regression [100, 102].